OpenStudy (anonymous):

Kyoko just tu ned 30 and gabe birth to a baby girl. She knows that when her daughter finishes high school, a post secondary educaton will cost much more than it does today. Kyokocplans to put $10 per week into her daughter's RESP (registered education savings plan). In addition, the federal government will contribute 20% of the investor's RESP contribution each year up to a maximum of $400 per year. How much will Kyoko have invested by her daughter's first birthday?

OpenStudy (anonymous):

First birthday = 52 weeks old. 52 weeks @ $10 per week =$520 contribution after a year. However, the federal government will contribute 20% of the investor's contribution, so $520*0.2=104. 104+520=624, total investment.

OpenStudy (anonymous):

The book said that the answer is $520

OpenStudy (anonymous):

my mistake. the question says how much will Kyoko have invested, so you ignore the government's contributions. $10 per week for 52 weeks (1 year) =$520

OpenStudy (anonymous):

52 x 10=520? Right?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

There is a follow up question! How much money will be in the fund after 1 year?

OpenStudy (anonymous):

Haha, that's what I answered in the first place! 52 weeks @ $10 per week =$520 contribution after a year. However, the federal government will contribute 20% of the investor's contribution, so $520*0.2=104. 104+520=624, total investment.